# Cap2Sum: Learning to Summarize Videos by Generating Captions

### v1

@@ 4 The TVSum-Caption and SumMe-Caption datasets
Since our method can be fine-tuned by the video captions, we need appropriate data (i.e., videos with both caption and summary annotation) to verify the generalization capacity of Cap2Sum under this weakly-supervised fine-tuning scheme. Thus, in this section, we propose the TVSum-Caption and SumMe-Caption datasets. These two datasets are based on the TVSum Song et al. (2015) and SumMe Gygli et al. (2014) datasets that are commonly used in video summarization research. Specifically, we label the dense video captions for the videos in these two datasets. To generate objective video captions, each video is annotated by 5 annotators, and an experienced supervisor combines the annotations and gives the final annotation. We count the duration and number of events of the videos in the proposed datasets and compare them with the AcitivityNet-Caption (ANet) dataset (i.e the training dataset employed in our experiments), the results are shown in Fig. 3. Through the comparison, we notice that compared to the videos in ANet, the videos in TVSum-Caption tend to be longer and contain more events. In addition, although the videos in SumMe-Caption have a similar length and number of events as the videos in ANet, SumMe-Caption contains a large number of first-person videos, which are rarely found in ANet. Therefore, the generalization from ANet to TVSum-Caption and SumMe-caption dataset is challenging. Thus, by conducting experiments on the proposed dataset, we can convincingly verify the generalization capacity of Cap2Sum. The TVSum-Caption and SumMe-Caption dataset will be publicly released to support further research on caption-based video summarization.
@@ 5.1 Settings
We pre-train Cap2Sum on the ActivityNet-Caption (ANet) dataset Wang et al. (2018). ANet is a large-scale dense video caption dataset that contains over 14000 annotated videos. The fine-tuning and evaluation are conducted on the video summarization datasets TVSum Song et al. (2015) and SumMe Gygli et al. (2014). TVSum dataset contains 50 videos, each video has frame-wise importance scores labelled by 20 annotators. SumMe dataset contains 25 videos, and also has frame-wise scores labelled by 15-18 annotators. The proposed caption annotations for TVSum and SumMe datasets are also employed to examine the generalization capacity of Cap2Sum under weakly-supervised fine-tuning.
@@
In the previous works Zhang et al. (2016b, a); Rochan et al. (2018), the evaluation of video summarization is usually conducted on the TVSum and SumMe datasets in three data settings: Standard, Augment, and Transfer. However, since Cap2Sum requires weakly-supervised pre-training on the large-scale video caption dataset, experiments using Standard setting (i.e., training and testing splits are from the same dataset) are not available. Thus, our experiments are conducted using Augment and Transfer settings. In the Augment setting, the model is pre-trained on the ANet dataset and fine-tuned on the SumMe or TVSum dataset by 80% of the videos following previous works. In the Transfer setting, the fine-tuning and testing are conducted on the different datasets (e.g., fine-tuning on the SumMe dataset and testing on the TVSum dataset). We also conduct experiments using zero-shot setting (i.e., without fine-tuning) to compare the performance with other unsupervised and weakly-supervised methods.
@@
We employ the pre-trained image encoder (ViT-B/16) of CLIP Radford et al. (2021) as the visual encoder in Cap2Sum. The threshold $\tau$ of CLIP Prior mechanism is set to 0.4. $\beta_{giou}$, $\beta_{cls}$, $\beta_{ec}$, and $\beta_{pred}$ are set to 4, 2, 0.5, and 0.5 respectively. $\beta_{cap}$, $\beta_{prior}$, $\beta_{len}$, and $\beta_{var}$ are set to 2, 10, 0.5, and 0.5 respectively. The model is trained by Adam optimizer Kingma and Ba (2014) with the learning rate set to 5e-5 and batch size set to 1. Following the previous works Zhang et al. (2016b, a); Rochan et al. (2018); Narasimhan et al. (2021), we report F1-Score to evaluate the performance of video summarization.
@@
SumTransfer Zhang et al. (2016a) CVPR 2016 41.3 38.5 - - 17
dppLSTM Zhang et al. (2016b) ECCV 2016 42.9 41.8 59.6 58.7 16
SUM-GANsup Mahasseni et al. (2017) CVPR 2017 43.6 - 61.2 - 14
SUM-FCN Rochan et al. (2018) ECCV 2018 51.1 44.1 59.2 58.2 11
SUM-DeepLab Rochan et al. (2018) ECCV 2018 50.2 45.0 59.1 57.4 12
DR-DSN Zhou et al. (2018a) AAAI 2018 43.9 42.6 59.8 58.9 13
$re$-SEQ2SEQ Zhang et al. (2018) ECCV 2018 44.9 - 63.9 - 9
VASNet Fajtl et al. (2018) ACCV 2018 51.1 - 62.4 - 8
UnpairedVSN Rochan and Wang (2019) CVPR 2019 48.0 41.6 56.1 55.7 15
SumGraph Park et al. (2020) ECCV 2020 52.9 48.7 65.8 60.5 6
RSGN Zhao et al. (2021) TPAMI 2021 45.7 44.0 61.1 60.0 10
CLIP-It Narasimhan et al. (2021) NeurIPS 2021 56.4 51.9 69.0 65.5 3
RR-STG Zhu et al. (2022) TIP 2022 54.8 45.4 63.6 59.7 7
iPTNet Jiang and Mu (2022) ECCV 2022 56.9 49.2 64.2 59.8 5
A2Summ He et al. (2023) CVPR 2023 - 55.0 - 63.4 4
@@ 5.2 Ablation Studies
In this subsection, we conduct experiments to verify the effectiveness of weakly-supervised fine-tuning (i.e., supervised by the dense video caption) on the proposed TVSum-Caption and SumMe-Caption datasets. We also conduct ablation studies on the training split ratio of supervised fine-tuning to verify the few-shot performance of Cap2Sum. We report the F1-Scores of unsupervised setting (unsup.), weakly-supervised setting (weak sup.), and supervised setting using $p$videos to fine-tune (sup. $p$videos) in Table 1. The results show that using video captions to fine-tune Cap2Sum is very effective (gains 2.4% improvement on SumMe and 4.2% improvement on TVSum compared to unsupervised setting). This means that the user only needs to provide part of the video captions of the videos that need to be summarized, and Cap2Sum can be effectively generalized to that user’s videos. In addition, the comparison of different training split ratios of supervised fine-tuning demonstrates Cap2Sum can achieve promising performance using a few videos to fine-tune. Thus, it is proved that Cap2Sum also has a strong few-shot video summarization performance.
@@
The proposed CLIP Prior mechanism is the key to reduce the domain gap and improve the generalization capacity. To verify the effectiveness of CLIP Prior mechanism, we compare the F1-Scores of removing our CLIP Prior mechanism during pre-training or fine-tuning (we employ weakly-supervised fine-tuning in this experiment) in Table 2. Obviously, introducing our CLIP Prior mechanism during pre-training significantly improves performance. And CLIP-Prior mechanism is also effective when introduced to fine-tuning. These results demonstrate that the proposed CLIP Prior mechanism is very effective in improving the generalization capacity of the pre-trained Cap2Sum model.
@@
SUM-FCNunsup Rochan et al. (2018) ECCV 2018 39.5 -
UnpairedVSN Rochan and Wang (2019) CVPR 2019 41.6 55.7
Cycle-SUM Li et al. (2019) CVPR 2019 41.9 57.6
ACGAN He et al. (2019) ACMMM 2019 44.5 57.8
SumGraph Park et al. (2020) ECCV 2020 47.0 57.6
ERA Wu et al. (2021) BMVC 2021 48.8 58.0
DSR-RL-GRU Phaphuangwittayakul et al. (2021) ICME 2021 50.3 60.2
CLIP-It Narasimhan et al. (2021) NeurIPS 2021 50.0 62.8
CA-SUM Apostolidis et al. (2022) ACM ICMR 2022 51.1 61.4
@@ 5.3 Comparison to Supervised Methods
In this section, we compare our approach to those state-of-the art supervised video summarization methods. The experiments are conducted using Augmented and Transfer setting, we report the F1-Scores on TVSum and SumMe datasets in Table 3. As the results show, Cap2Sum fine-tuned by ground-truth summarization (i.e., Cap2Sumsup) outperforms the state-of-the-art baselines on all settings and on both datasets, thus demonstrating that pre-training using text captions on large-scale datasets is effective in improving the performance. Particularly, compared to CLIP-It Narasimhan et al. (2021), which is also a caption-based method, our method achieves significant performance improvement, proving that our approach exploits the video captions in a more efficient way. We also compare the performance of Cap2Sum fine-tuned by video captions (i.e., Cap2Sumwsup) and the results show that Cap2Sum also achieves very competitive performance using this setting. Therefore, the effectiveness of learning to summarize videos by generating captions is proven.
@@ 5.4 Comparison to Unsupervised Methods
Since Cap2Sum is able to be trained without ground-truth summarization, but by other weak labels, similar to the previous unsupervised video summarization methods Khosla et al. (2013); Mahasseni et al. (2017). It is fair to compare our method (without fine-tuning) with other unsupervised methods. We use the Transfer setting for fair comparison because the videos in the testing dataset are not introduced during the training phase of unsupervised Cap2Sum (Cap2Sumusup). We report the F1-Scores in Table 4. In this zero-shot scenario, Cap2Sum outperforms the previous methods on both datasets. We again observe significant performance improvements compared to CLIP-It Narasimhan et al. (2021) and our Cap2Sum stands state-of-the-art performance among all approaches, which strongly demonstrates the effectiveness of our approach.
@@ 5.5 Visualization
Fig. 4 shows qualitative results with the captions generated by Cap2Sum on TVSum. We pick 4 video clips with the highest caption confidence and proposal score in each video. We observe that the generated captions have a good consistency with the corresponding video segments and ground-truth importance scores, enabling to train a robust and accurate video summarizer.

### v2

@@ 4 The TVSum-Caption and SumMe-Caption datasets
Since our method can be fine-tuned by the video captions, we need appropriate data (i.e., videos with both caption and summary annotation) to verify the generalization capacity of Cap2Sum under this weakly-supervised fine-tuning scheme. Thus, in this section, we propose the TVSum-Caption and SumMe-Caption datasets. These two datasets are based on the TVSum [29] and SumMe [5] datasets that are commonly used in video summarization research. Specifically, we label the dense video captions for the videos in these two datasets. To generate objective video captions, each video is annotated by 5 annotators, and an experienced supervisor combines the annotations and gives the final annotation. We count the duration and number of events of the videos in the proposed datasets and compare them with the AcitivityNet-Caption (ANet) dataset (i.e the training dataset employed in our experiments), the results are shown in Fig.˜3. Through the comparison, we notice that compared to the videos in ANet, the videos in TVSum-Caption tend to be longer and contain more events. In addition, although the videos in SumMe-Caption have a similar length and number of events as the videos in ANet, SumMe-Caption contains a large number of first-person videos, which are rarely found in ANet. Therefore, the generalization from ANet to TVSum-Caption and SumMe-caption dataset is challenging. Thus, by conducting experiments on the proposed dataset, we can convincingly verify the generalization capacity of Cap2Sum. The TVSum-Caption and SumMe-Caption dataset will be publicly released to support further research on caption-based video summarization.
@@ 5.1 Settings
We pre-train Cap2Sum on the ActivityNet-Caption (ANet) dataset [35]. ANet is a large-scale dense video caption dataset that contains over 14000 annotated videos. The fine-tuning and evaluation are conducted on the video summarization datasets TVSum [29] and SumMe [5]. TVSum dataset contains 50 videos, each video has frame-wise importance scores labelled by 20 annotators. SumMe dataset contains 25 videos, and also has frame-wise scores labelled by 15-18 annotators. The proposed caption annotations for TVSum and SumMe datasets are also employed to examine the generalization capacity of Cap2Sum under weakly-supervised fine-tuning.
@@
In the previous works [41, 40, 28], the evaluation of video summarization is usually conducted on the TVSum and SumMe datasets in three data settings: Standard, Augment, and Transfer. However, since Cap2Sum requires weakly-supervised pre-training on the large-scale video caption dataset, experiments using Standard setting (i.e., training and testing splits are from the same dataset) are not available. Thus, our experiments are conducted using Augment and Transfer settings. In the Augment setting, the model is pre-trained on the ANet dataset and fine-tuned on the SumMe or TVSum dataset by 80% of the videos following previous works. In the Transfer setting, the fine-tuning and testing are conducted on the different datasets (e.g., fine-tuning on the SumMe dataset and testing on the TVSum dataset). We also conduct experiments using zero-shot setting (i.e., without fine-tuning) to compare the performance with other unsupervised and weakly-supervised methods.
@@
We employ the pre-trained image encoder (ViT-B/16) of CLIP [25] as the visual encoder in Cap2Sum. The threshold $\tau$ of CLIP Prior mechanism is set to 0.4. $\beta_{giou}$, $\beta_{cls}$, $\beta_{ec}$, and $\beta_{pred}$ are set to 4, 2, 0.5, and 0.5 respectively. $\beta_{cap}$, $\beta_{prior}$, $\beta_{len}$, and $\beta_{var}$ are set to 2, 10, 0.5, and 0.5 respectively. The model is trained by Adam optimizer [15] with the learning rate set to 5e-5 and batch size set to 1. Following the previous works [41, 40, 28, 21], we report F1-Score to evaluate the performance of video summarization.
@@
SumTransfer [40] CVPR 2016 41.3 38.5 - - 17
dppLSTM [41] ECCV 2016 42.9 41.8 59.6 58.7 16
SUM-GANsup [20] CVPR 2017 43.6 - 61.2 - 14
SUM-FCN [28] ECCV 2018 51.1 44.1 59.2 58.2 11
SUM-DeepLab [28] ECCV 2018 50.2 45.0 59.1 57.4 12
DR-DSN [45] AAAI 2018 43.9 42.6 59.8 58.9 13
rere-SEQ2SEQ [42] ECCV 2018 44.9 - 63.9 - 9
VASNet [4] ACCV 2018 51.1 - 62.4 - 8
UnpairedVSN [27] CVPR 2019 48.0 41.6 56.1 55.7 15
SumGraph [23] ECCV 2020 52.9 48.7 65.8 60.5 6
RSGN [44] TPAMI 2021 45.7 44.0 61.1 60.0 10
CLIP-It [21] NeurIPS 2021 56.4 51.9 69.0 65.5 3
$R$-STG [47] TIP 2022 54.8 45.4 63.6 59.7 7
iPTNet [11] ECCV 2022 56.9 49.2 64.2 59.8 5
A2Summ [7] CVPR 2023 - 55.0 - 63.4 4
@@ 5.2 Ablation Studies
In this subsection, we conduct experiments to verify the effectiveness of weakly-supervised fine-tuning (i.e., supervised by the dense video caption) on the proposed TVSum-Caption and SumMe-Caption datasets. We also conduct ablation studies on the training split ratio of supervised fine-tuning to verify the few-shot performance of Cap2Sum. We report the F1-Scores of unsupervised setting (unsup.), weakly-supervised setting (weak sup.), and supervised setting using pp% videos to fine-tune (sup. pp% videos) in Table˜1. The results show that using video captions to fine-tune Cap2Sum is very effective (gains 2.4% improvement on SumMe and 4.2% improvement on TVSum compared to unsupervised setting). This means that the user only needs to provide part of the video captions of the videos that need to be summarized, and Cap2Sum can be effectively generalized to that user’s videos. In addition, the comparison of different training split ratios of supervised fine-tuning demonstrates Cap2Sum can achieve promising performance using a few videos to fine-tune. Thus, it is proved that Cap2Sum also has a strong few-shot video summarization performance.
@@
The proposed CLIP Prior mechanism is the key to reduce the domain gap and improve the generalization capacity. To verify the effectiveness of CLIP Prior mechanism, we compare the F1-Scores of removing our CLIP Prior mechanism during pre-training or fine-tuning (we employ weakly-supervised fine-tuning in this experiment) in Table˜2. Obviously, introducing our CLIP Prior mechanism during pre-training significantly improves performance. And CLIP-Prior mechanism is also effective when introduced to fine-tuning. These results demonstrate that the proposed CLIP Prior mechanism is very effective in improving the generalization capacity of the pre-trained Cap2Sum model.
@@
SUM-FCNunsup [28] ECCV 2018 39.5 -
UnpairedVSN [27] CVPR 2019 41.6 55.7
Cycle-SUM [18] CVPR 2019 41.9 57.6
ACGAN [8] ACMMM 2019 44.5 57.8
SumGraph [23] ECCV 2020 47.0 57.6
ERA [37] BMVC 2021 48.8 58.0
DSR-RL-GRU [24] ICME 2021 50.3 60.2
CLIP-It [21] NeurIPS 2021 50.0 62.8
CA-SUM [2] ACM ICMR 2022 51.1 61.4
@@ 5.3 Comparison to Supervised Methods
In this section, we compare our approach to those state-of-the art supervised video summarization methods. The experiments are conducted using Augmented and Transfer setting, we report the F1-Scores on TVSum and SumMe datasets in Table˜3. As the results show, Cap2Sum fine-tuned by ground-truth summarization (i.e., Cap2Sumsup) outperforms the state-of-the-art baselines on all settings and on both datasets, thus demonstrating that pre-training using text captions on large-scale datasets is effective in improving the performance. Particularly, compared to CLIP-It [21], which is also a caption-based method, our method achieves significant performance improvement, proving that our approach exploits the video captions in a more efficient way. We also compare the performance of Cap2Sum fine-tuned by video captions (i.e., Cap2Sumwsup) and the results show that Cap2Sum also achieves very competitive performance using this setting. Therefore, the effectiveness of learning to summarize videos by generating captions is proven.
@@ 5.4 Comparison to Unsupervised Methods
Since Cap2Sum is able to be trained without ground-truth summarization, but by other weak labels, similar to the previous unsupervised video summarization methods [14, 20]. It is fair to compare our method (without fine-tuning) with other unsupervised methods. We use the Transfer setting for fair comparison because the videos in the testing dataset are not introduced during the training phase of unsupervised Cap2Sum (Cap2Sumusup). We report the F1-Scores in Table˜4. In this zero-shot scenario, Cap2Sum outperforms the previous methods on both datasets. We again observe significant performance improvements compared to CLIP-It [21] and our Cap2Sum stands state-of-the-art performance among all approaches, which strongly demonstrates the effectiveness of our approach.
@@ 5.5 Visualization
Fig.˜4 shows qualitative results with the captions generated by Cap2Sum on TVSum. We pick 4 video clips with the highest caption confidence and proposal score in each video. We observe that the generated captions have a good consistency with the corresponding video segments and ground-truth importance scores, enabling to train a robust and accurate video summarizer.